The length of a rectangle is 7 cm longer than the width.if the perimeter is 62 cm,find the length.
Please help me!
Answer by jim_thompson5910(35256) 
 The length is 7 cm longer than the width, so L = W+7. We'll use the perimeter formula P = 2*(L+W) where the given perimeter is P = 62
P = 2*(L+W)
P = 2*(W+7+W) ... plug in L = W+7
62 = 2*(2W+7) ... plug in P = 62. Solve for W
62 = 4W + 14
62-14 = 4W
48 = 4W
4W = 48
W = 48/4
W = 12
Since W = 12, we can use this to find L
L = W+7
L = 12+7
L = 19
So the length is 19 cm.
------------------------------------------------------------------------------------------------------------------------
If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).
Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.
Jim
------------------------------------------------------------------------------------------------------------------------
|
|
|